### v1

# Lowering the Error Floor of Error Correction Code Transformer

Abstract
With the success of transformer architectures across diverse applications, the error correction code transformer (ECCT) has gained significant attention for its superior decoding performance. In spite of its advantages, the error floor phenomenon in ECCT decoding remains unexplored. We present the first investigation of the error floor issue in ECCT and propose a hybrid decoding approach that integrates hard decision decoders as pre- and post-decoders with ECCT to effectively lower the error floor. In particular, we introduce a novel loss function for ECCT that considers the dynamics of hybrid decoding algorithm. Training ECCT with the proposed loss function enhances its ability to correct specific error patterns by taking into account its interaction with the auxiliary decoders. Simulation results demonstrate that the proposed hybrid decoder with the novel loss function significantly outperforms the original ECCT in both the waterfall and the error floor regions.
I Introduction
The significant success of the transformer architecture [1] across various applications [2, 3] has inspired the development of a transformer-based error correction code decoder, known as the error correction code transformer (ECCT) [4], which has achieved notable decoding performance, particularly for short codes. ECCT effectively captures the relevance between bits using the attention mechanism, while leveraging the code structure by masking the attention map based on the parity check matrix. In line with this approach, Choukroun and Wolf presented a foundation model for ECCT [5] and the end-to-end optimization of both the encoder and decoder [6]. Furthermore, Park et al. proposed novel transformer architectures that utilize the multiple mask matrices [7] and the cross-attention mechanism [8]. These transformer-based decoders [4, 5, 6, 7, 8] have shown superior decoding performance compared to conventional decoding algorithms and other neural decoders.
In spite of its notable decoding performance in the waterfall region, the error floor of ECCT has yet to be investigated. The error floor refers to the phenomenon where the decrease in error probability slows significantly in the high signal-to-noise ratio (SNR) region, compared to the waterfall region [9]. For emerging applications demanding extremely low error rate, such as the next-generation ultra-reliable and low-latency communications (xURLLC) [10], it is important to mitigate the error floor problem [11]. In particular, several techniques have been developed to mitigate the error floor phenomenon in low-density parity-check (LDPC) codes by leveraging decoder diversity [12] and boosting-based learning [11]. However, to the best of our knowledge, the error floor in transformer-based decoders remains unexplored.
In this paper, we present the first investigation into the error floor phenomenon in transformer-based decoders, which has neither been observed nor discussed in prior studies. Specifically, we find that ECCT suffers from an error floor even when decoding Bose-Chaudhuri-Hocquenghem (BCH) codes. Our simulation results reveal that this error floor prevents ECCT from achieving extremely low frame error rates (FERs) in the high SNR region. This finding motivates the development of decoding algorithms to mitigate the error floor in ECCT.
To address this issue, we propose a simple yet effective hybrid decoding approach that effectively lowers the error floor in ECCT decoding. The proposed hybrid decoder integrates hard decision decoders as pre- and post-decoders for ECCT. This approach is motivated by the observation that ECCT can reduce errors in received vectors, even when it fails to accurately estimate the transmitted codewords. By leveraging hard decision decoders, the hybrid decoder can correct a broader range of error patterns, significantly improving decoding performance in both the error floor and waterfall regions.
(a) BCH codes
(b) LDPC codes
Figure 1: Error floor phenomenon of ECCT: (a) FER performance of ECCT for BCH codes, and (b) FER performance of ECCT for LDPC codes.
In particular, we propose a novel loss function for the hybrid decoder to further improve its decoding performance. This loss function is specifically designed to enhance the interactions between ECCT and the pre- and post-hard decision decoders. First, considering the pre-decoder, the training dataset is constrained to include only the received vectors with errors exceeding the error correction capability of the hard decision decoder. Second, for the post-decoder, the loss function is adjusted depending on whether the number of errors in ECCT’s output exceeds the error correction capability. Rather than training ECCT to correct all errors for all received vectors as in [4], the proposed loss function enables ECCT to focus on reducing critical errors in specific received vectors. This targeted approach enhances error correction throughout the overall hybrid decoding process.
II Preliminaries
II-A Notation
We denote an $(n,k)$ linear block code as $\mathcal{C}$, which is characterized by a generator matrix $G$, a parity check matrix $H$, and a minimum Hamming distance $d_{\text{min}}$. The message and the codeword are denoted by $\mathbf{m}\in\{0,1\}^{k}$ and $\mathbf{x}\in\mathcal{C}\subseteq\{0,1\}^{n}$. Additionally, $\mathbf{X}$ represents the random variable of $\mathbf{x}$. We consider the binary phase shift keying (BPSK) modulation with the additive white Gaussian noise (AWGN) channel. The modulated signal and the received vector are denoted by $\mathbf{x}_{s}$ and $\mathbf{y}$ with the noise vector $\mathbf{z}\sim\mathcal{N}(0,\sigma^{2}I)$. The Hamming distance between two codewords $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$ is denoted by $d_{H}(\mathbf{x}_{1},\mathbf{x}_{2})$. The decoder $f(\cdot)$ outputs the estimated codeword $\widehat{\mathbf{x}}$ from the received vector $\mathbf{y}$. The hard decision decoder corrects all the received vectors $\mathbf{y}$ such that $d_{H}(\mathbf{x},\mathbf{y})\leq t_{c}$, where the error correction capability is denoted as $t_{c}=\lfloor(d_{\text{min}}-1)/2\rfloor$.
II-B Error Correction Code Transformer
ECCT is a model-free neural decoder based on the transformer architecture [4]. As in [13], ECCT preprocesses the input as $h(\mathbf{y})=[|\mathbf{y}|,s(\mathbf{y})]$, where [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes the concatenation of two vectors and $s(\mathbf{y})$ represents the syndrome $H\mathbf{y}_{b}=H\text{bin}(\text{sign}(\mathbf{y}))$. Here, $\text{sign}(\cdot)$ is the sign function such that $\text{sign}(\alpha)=1$ if $\alpha\geq 0$ and $-1$ otherwise; and $\text{bin}(\cdot)$ is the binarization function defined as $\text{bin}(\alpha)=0.5(1-\alpha)$. Each element of the preprocessed input is then linearly mapped into an embedding vector of dimension $d$ for the attention layers.
The code-aware self-attention operation is defined as [4]:
$A_{H}\left(Q,K,V\right)=\text{Softmax}(d^{-1/2}(QK^{\mathsf{T}}+{g}(H)))V,$
where $Q$, $K$, $V$, and ${g}(H)$ denote the query, key, value, and mask matrix, respectively. The mask matrix $g(H)$ is constructed using the parity check matrix $H$, enabling the self-attention to focus selectively on the important relations between the embedding vectors.
The decoder of ECCT, denoted as $f_{\boldsymbol{\theta}}(\mathbf{y})$ with trainable parameters $\boldsymbol{\theta}$, aims to predict the multiplicative noise $\tilde{\mathbf{z}}_{s}$ [4]. The multiplicative noise is defined by $\mathbf{y}=\mathbf{x}_{s}+\mathbf{z}=\mathbf{x}_{s}\odot\tilde{\mathbf{z}}_{s}$, where $\odot$ denotes element-wise multiplication. ECCT is trained using the following binary cross entropy loss function:
$l_{\text{BCE}}(\mathbf{x},f_{\boldsymbol{\theta}}(\mathbf{y}))$ $=-\sum_{i=1}^{n}\{\tilde{z}_{i}\log(1-\sigma(f_{\boldsymbol{\theta}}(\mathbf{y})_{i}))$
$\quad+(1-\tilde{z}_{i})\log\sigma(f_{\boldsymbol{\theta}}(\mathbf{y})_{i})\},$ (1)
where $\tilde{\mathbf{z}}=\text{bin}(\text{sign}(\tilde{\mathbf{z}}_{s}))$ denotes the binarized multiplicative noise and $\sigma(\cdot)$ denotes the sigmoid function $\sigma(\alpha)=1/(1+e^{-\alpha})$. Lastly, the estimated codeword can be obtained as
$\widetilde{f}_{\boldsymbol{\theta}}(\mathbf{y})=\text{bin}(\text{sign}(f_{\boldsymbol{\theta}}(\mathbf{y})\odot\mathbf{y})).$ (2)
III Error Floor of Error Correction Code Transformer and Hybrid Decoders
We report the error floor phenomenon of ECCT for the first time and propose a hybrid decoding approach to address the error floor. Fig. 1 shows the error floor phenomenon of ECCT when decoding BCH and LDPC codes. ECCT suffers from the error floor for both BCH and LDPC codes, preventing it from achieving extremely low FER in the high SNR region. For BCH codes, it is unexpected that ECCT exhibits an error floor, as this phenomenon does not occur in conventional decoders. In addition, the results for LDPC codes imply that ECCT cannot avoid the error floor, which is a critical issue even in the belief propagation (BP) algorithm.
(a) Hybrid decoding with hard decision pre- and post-decoders
(b) Hybrid decoding with hard decision pre-decoder
(c) Hybrid decoding with hard decision post-decoder
Figure 2: Block diagrams of hybrid decoding architectures.
We propose a hybrid decoding approach that combines ECCT with hard decision decoders to lower the error floor of ECCT. The block diagrams of the proposed architectures are shown in Fig. 2. As shown in Figs. 2LABEL:sub@fig:architecture_pre_post, LABEL:sub@fig:architecture_pre, and LABEL:sub@fig:architecture_post, hard decision decoders can be employed as both pre- and post-decoder, or solely as a pre-decoder or a post-decoder, respectively. Since hard decision decoders and ECCT can correct different classes of error patterns, the hybrid decoding approach can correct a broader range of errors, effectively lowering the error floor.
In the hybrid decoding architecture with hard decision pre- and post-decoders (see Figs. 2LABEL:sub@fig:architecture_pre_post), the decoding process terminates if the hard decision pre-decoder successfully estimates the correct codewords; otherwise, it proceeds to the next stage. If the pre-decoder fails, the received vector $\mathbf{y}$ is passed to ECCT for decoding in the subsequent stage. If ECCT also fails, the process continues to the hard decision post-decoder. Unlike the pre-decoder, which operates on the received vector $\mathbf{y}$, the post-decoder takes the output of ECCT as its input. The ECCT output typically contains fewer errors since ECCT can reduce errors even when it fails to decode. Depending on the situation, either the pre-decoder or the post-decoder can be omitted.
(a) $d_{H}(\mathbf{x},\mathbf{y})\leq t_{c}$
(b) $d_{H}(\mathbf{x},\mathbf{y})>t_{c}$
Figure 3: Histograms of the number of errors in the output vector of ECCT in decoding (63, 45) BCH codes.
The roles of the pre-decoder and post-decoder in hybrid decoding can be better understood by examining the distribution of errors in the output of ECCT. Fig. 3 shows the histogram of the number of errors in the output vectors of ECCT for the $(63,45)$ BCH code. Figs. 3LABEL:sub@fig:error_distribution-a and LABEL:sub@fig:error_distribution-b show the histogram of received vectors with errors less than or equal to $t_{c}$ and the histogram of the received vectors with errors more than $t_{c}$, respectively. As shown in Fig. 3LABEL:sub@fig:error_distribution-a, ECCT may fail to correct all errors, in some cases, even increase the number of errors for received vectors that initially have $t_{c}$ or fewer errors. Since the hard decision decoder can correct all the received vectors with $t_{c}$ or fewer errors, using it as a pre-decoder improves the overall decoding performance. For some received vectors with more than $t_{c}$ errors, ECCT is capable of estimating the correct codewords. Even when it fails, it frequently reduces the number errors as shown in Fig. 3LABEL:sub@fig:error_distribution-b. This behavior occurs since the binary cross entropy loss (II-B) trains ECCT to reduce errors in its output. Consequently, employing a hard decision post-decoder can correct the remaining errors ($\leq t_{c}$), thereby enhancing the overall decoding performance.
IV Proposed Loss Function for Hybrid Decoders
This section proposes a training method for the hybrid decoding algorithm. To improve the error correction performance of the hybrid decoder, the ECCT decoder should be trained with consideration for its integration with hard decision decoders. Specifically, we focus on the hybrid decoding scheme that combines ECCT with both pre- and post-decoders, as shown in Fig. 2LABEL:sub@fig:architecture_pre_post. Suppose that $\widehat{\mathbf{x}}_{1}=f_{\text{pre}}(\mathbf{y})$, $\widehat{\mathbf{x}}_{2}=f_{\boldsymbol{\theta}}(\widehat{\mathbf{x}}_{1})$, and $\widehat{\mathbf{x}}=f_{\text{post}}(\widehat{\mathbf{x}}_{2})$ represent the outputs of the pre-decoder, ECCT, and the post-decoder, respectively. The bit error rate (BER) loss for this hybrid decoder is given by
$\mathcal{L}(\boldsymbol{\theta})$ $=\mathbb{E}_{\mathbf{X},\widehat{\mathbf{X}}}\left[d_{H}(\mathbf{X},\widehat{\mathbf{X}})\right]$
$=\mathbb{E}_{\mathbf{X},\mathbf{Y}}\left[d_{H}(\mathbf{X},f_{\text{post}}\circ f_{\boldsymbol{\theta}}\circ f_{\text{pre}}(\mathbf{Y}))\right].$ (3)
However, ECCT cannot be trained directly using the loss (3) since the hard decision decoders $f_{\text{pre}}$ and $f_{\text{pre}}$ are not differentiable. To enable gradient-based training of ECCT, we derive the following equivalent loss for hybrid decoding:
Lemma 1
Consider the hybrid decoding algorithm with the hard decision decoders as pre- and post-decoders. The loss (3) is equivalent to:
$\mathbb{E}_{\mathbf{X},\mathbf{Y}}\Big{[}$ $u(d_{H}(\mathbf{X},\widetilde{f}_{\boldsymbol{\theta}}(\mathbf{Y}))-t_{c})$
×$\times d_{H}(\mathbf{X},\widetilde{f}_{\boldsymbol{\theta}}(\mathbf{Y}))|d_{H}(\mathbf{X},\mathbf{Y})>t_{c}\Big{]},$ (4)
where $u(\alpha)$ denotes the step function, defined as $u(\alpha)=0$ if $\alpha\leq 0$ and $1$ otherwise.
Proof:
The expectation in (3) can be evaluated by considering two events: $d_{H}(\mathbf{x},\mathbf{y})\leq t_{c}$ and $d_{H}(\mathbf{x},\mathbf{y})>t_{c}$. Using the law of total expectation, we can expand the loss as follows:
$\mathcal{L}(\boldsymbol{\theta})$ $=\mathbb{E}_{\mathbf{X},\mathbf{Y}}\left[d_{H}(\mathbf{X},f_{\text{post}}\circ f_{\boldsymbol{\theta}}\circ f_{\text{pre}}(\mathbf{Y}))|d_{H}(\mathbf{X},\mathbf{Y})>t_{c}\right]$
×$\quad\times\Pr(d_{H}(\mathbf{X},\mathbf{Y})>t_{c})$ (5)
$=\mathbb{E}_{\mathbf{X},\mathbf{Y}}\left[d_{H}(\mathbf{X},f_{\text{post}}\circ f_{\boldsymbol{\theta}}(\mathbf{Y}))|d_{H}(\mathbf{X},\mathbf{Y})>t_{c}\right]$
×$\quad\times\Pr(d_{H}(\mathbf{X},\mathbf{Y})>t_{c}),$ (6)
where (5) holds since the hard decision pre-decoder can correct all errors whenever $d_{H}(\mathbf{x},\mathbf{y})\leq t_{c}$, and  (6) holds because the pre-decoder passes the received vector to the subsequent stage if it fails to decode.
Since the decoding success of the post-decoder is fully determined by the number of errors in the output vector of ECCT, the loss can be expressed as follows:
$\mathbb{E}_{\mathbf{X},\mathbf{Y}}\left[u(d_{H}(\mathbf{X},\widetilde{f}_{\boldsymbol{\theta}}(\mathbf{Y}))-t_{c})d_{H}(\mathbf{X},\widetilde{f}_{\boldsymbol{\theta}}(\mathbf{Y}))|d_{H}(\mathbf{X},\mathbf{Y})>t_{c}\right]$
×$\times\Pr(d_{H}(\mathbf{X},\mathbf{Y})>t_{c}).$
By omitting $\Pr(d_{H}(\mathbf{X},\mathbf{Y})>t_{c})$, which is independent of $\boldsymbol{\theta}$, we obtain (4). ∎
(a) (31, 16) BCH code
(b) (63, 36) BCH code
(c) (63, 45) BCH code
(d) (63, 51) BCH code
Figure 4: FER performance of ML decoding [14], hard decision decoding, ECCT [4], and hybrid decoding.
In Lemma 1, the pre-decoder is addressed through a conditional expectation, ensuring that ECCT is trained exclusively on received vectors $\mathbf{y}$ containing more than $t_{c}$ errors, which the pre-decoder is unable to correct. In addition, the post-decoder is incorporated by multiplying the loss function with the step function $u(d_{H}(\mathbf{X},f_{\boldsymbol{\theta}}(\mathbf{Y}))-t_{c})$. ECCT trained with this loss aims to reduce errors to $t_{c}$ or fewer, rather than correcting all errors. The remaining errors, which are less than or equal to $t_{c}$, are corrected by the post-decoder. Furthermore, the loss (3) can be upper bounded as in the following proposition.
Proposition 1
The loss (3) can be upper bounded by
[…]
×$\times l_{\text{BCE}}(\mathbf{X},f_{\boldsymbol{\theta}}(\mathbf{Y}))|d_{H}(\mathbf{X},\mathbf{Y})>t_{c}\Big{]}.$ (7)
Proof:
Following Lemma 1, the loss $\mathcal{L}(\boldsymbol{\theta})$ can be reformulated as (4). In (4), $d_{H}(\mathbf{x},\widetilde{f}_{\boldsymbol{\theta}}(\mathbf{y}))=\sum_{i=1}^{n}d_{H}(x_{i},\widetilde{f}_{\boldsymbol{\theta}}(\mathbf{y})_{i})$ is equivalent to $\sum_{i=1}^{n}l_{0\text{--}1}((\tilde{z}_{s})_{i}\cdot f_{\boldsymbol{\theta}}(\mathbf{y})_{i})$, where $l_{0\text{--}1}(\alpha)$ is defined as $l_{0\text{--}1}(\alpha)=1$ if $\alpha\leq 0$ and 00 otherwise. Here, $l_{0\text{--}1}(\alpha)$ can be upper bounded by the logistic loss function as follows [15]:
$l_{0\text{--}1}((\tilde{z}_{s})_{i}\cdot f_{\boldsymbol{\theta}}(\mathbf{y})_{i})\leq-\log_{2}(\sigma((\tilde{z}_{s})_{i}\cdot f_{\boldsymbol{\theta}}(\mathbf{y})_{i})).$
Since this logistic loss is equivalent to the scaled binary cross entropy loss in (II-B), the loss function (4) can be upper bounded by (7). ∎
In practice, ~$\widetilde{\mathcal{L}}({\boldsymbol{\theta}})$ can be computed using the following sample average, omitting the constant scaling term:
~$\widetilde{\mathcal{L}}({\boldsymbol{\theta}})$ $\simeq\frac{1}{M}\sum_{i=1}^{M}\Big{\{}u\big{(}d_{H}(\mathbf{x}^{(i)},\widetilde{f}_{\boldsymbol{\theta}}(\mathbf{y}^{(i)}))-t_{c}\big{)}$
×$\quad\times l_{\text{BCE}}\big{(}\mathbf{x}^{(i)},f_{\boldsymbol{\theta}}(\mathbf{y}^{(i)})\big{)}\Big{\}},$ (8)
where $\mathbf{x}^{(i)}$ can be fixed as the zero codeword and $\mathbf{y}^{(i)}$ is sampled as $\mathbf{x}_{s}^{(i)}+\mathbf{z}^{(i)}$, retaining only the samples that satisfy $d_{H}(\mathbf{x}^{(i)},\mathbf{y}^{(i)})>t_{c}$. The noise variance $\sigma^{2}$ is uniformly sampled from the training SNR range. In addition, the step function $u(x)$ is implemented via a straight-through estimator (STE) [16] with the sigmoid function as a proxy, as described in [17]. The Hamming distance $d_{H}(\mathbf{x},\widetilde{f}_{\boldsymbol{\theta}}(\mathbf{y}))$ is estimated by multiplying the code length $n$ with the soft BER loss defined in [18].
V Experimental Results
In this section, we present the experimental results for the proposed hybrid decoding algorithms. To evaluate its effectiveness, we measure the BER and FER performance for four BCH codes: $(31,16)$, $(63,36)$, $(63,45)$, and $(63,51)$, which have error correcting capabilities of $t=3$, $t=5$, $t=3$, and $t=2$, respectively. In all simulations involving ECCT, the number of decoding layers $N$ is fixed to $6$, and the embedding dimension $d$ is set to $128$. Additionally, the parity check matrices used in the experiments and maximum likelihood (ML) decoding results are obtained from [14].
Fig. 4 shows the FER performance of the proposed hybrid decoder along with other decoders: ML, hard decision, and the original ECCT. The hybrid decoder incorporates hard decision decoders as both pre- and post-decoders and is trained using the proposed loss (IV). The results indicate that the hybrid decoder effectively mitigates the error floor, significantly outperforming the original ECCT in the high SNR region. Furthermore, even in the waterfall region, the hybrid decoder outperforms both the original ECCT and hard decision decoder. Notably, the hybrid decoder achieves the ML decoding performance for the $(31,16)$ BCH code. By combining two distinct decoders and training them with the proposed loss, the hybrid decoding approach can achieve significant coding gains.
TABLE I: Comparison of FER of Decoders with Different Configurations
Methods (31, 16) BCH code (63, 45) BCH code
[…]
ECCT [4] $1.52\mathrm{e}{-2}$ $2.53\mathrm{e}{-3}$ $2.86\mathrm{e}{-4}$ $6.41\mathrm{e}{-2}$ $8.65\mathrm{e}{-3}$ $4.68\mathrm{e}{-4}$
Pre + ECCT $1.45\mathrm{e}{-2}$ $2.43\mathrm{e}{-3}$ $2.76\mathrm{e}{-4}$ $4.96\mathrm{e}{-2}$ $5.63\mathrm{e}{-3}$ $2.77\mathrm{e}{-4}$
ECCT + Post $6.70\mathrm{e}{-3}$ $9.23\mathrm{e}{-4}$ $7.47\mathrm{e}{-5}$ $3.36\mathrm{e}{-2}$ $3.42\mathrm{e}{-3}$ $1.26\mathrm{e}{-4}$
Pre + ECCT + Post $6.47\mathrm{e}{-3}$ $8.96\mathrm{e}{-4}$ $7.25\mathrm{e}{-5}$ $2.91\mathrm{e}{-2}$ $2.55\mathrm{e}{-3}$ $8.63\mathrm{e}{-5}$
Pre + ECCT + Post + loss (IV) $\mathbf{2.23\mathrm{\mathbf{e}}{-3}}$ $\mathbf{1.63\mathrm{\mathbf{e}}{-4}}$ $\mathbf{5.41\mathrm{\mathbf{e}}{-6}}$ $\mathbf{2.26\mathrm{\mathbf{e}}{-2}}$ $\mathbf{1.56\mathrm{\mathbf{e}}{-3}}$ $\mathbf{4.26\mathrm{\mathbf{e}}{-5}}$
To assess the impact of each contribution within the hybrid decoding framework, we compare the FER performance of different decoder configurations, as shown in Table I. In this table, “Pre” and “Post” represent hard decision pre- and post-decoders, respectively while “ECCT” refers to the original ECCT trained on the binary cross entropy loss function (II-B). The results show that the hybrid decoder combining ECCT with both pre- and post-decoders outperforms decoder configurations using only a single auxiliary decoder, implying the importance of both pre- and post-decoders. Moreover, the hybrid decoder trained with the proposed loss (IV) achieves the best FER performance among all configurations. This confirms that the proposed loss function effectively improves ECCT within the hybrid decoding framework, thereby enhancing overall decoding performance.
VI Conclusion
In this paper, we presented the first observation of the error floor problem of ECCT. To address this issue, we proposed hybrid decoding algorithms for ECCT, along with an optimized loss function. By integrating ECCT with hard decision decoders, the hybrid decoding approach effectively corrects a wider range of errors, thereby enhancing decoding performance. The proposed loss function enables ECCT to focus on correcting only critical errors, optimizing the overall hybrid decoding process. Our simulation results show that the hybrid decoder effectively lowers the error floor, achieving improved FER performance in the high SNR region. Moreover, the hybrid decoder significantly enhances decoding performance in the waterfall region. The proposed hybrid decoding algorithm and its associated training method, which accounts for the pre- and post-decoders, can be generalized to other neural decoders. This approach provides a versatile strategy for integrating neural decoders and conventional decoders to effectively lower the error floor and enhance decoding performance.
Acknowledgment
This work was partly supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (RS-2024-00398449, Network Research Center: Advanced Channel Coding and Channel Estimation Technologies for Wireless Communication Evolution) and the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2023-00212103).